OpenStudy (anonymous):

A hemisphere has a volume of 18π cm3. Find its radius

OpenStudy (dumbcow):

Volume is half that of a sphere V=(1/2)*(4/3)*pi*r^3 = 18pi -> 2/3 r^3 = 18 ->r^3 = 18*(3/2) -> r^3 = 27 = 3*3*3 r=3

OpenStudy (anonymous):

A sphere has a volume of 972π in.3. Find its radius.

OpenStudy (dumbcow):

same as above V = 4/3 pi r^3 = 972 solve for r

OpenStudy (anonymous):

A triangular pyramid has a volume of 180 cm3 and a height of 12 cm. Find the length of a side of the triangular base if the triangle’s height from that side is 6 cm.

OpenStudy (dumbcow):

volume of pyramid is area of base(B)*height V = B*12 = 180 B = 180/12 = 15 Area of base is area of triangle A = 1/2 * length(L)*height = 15 1/2*6*L = 15 solve for L

OpenStudy (dumbcow):

oops i messed up volume of pyramid is 1/3* B*h correction 1/3*B*12 = 180 B = 180/4 = 45 ... A=1/2*6*L = 45

OpenStudy (anonymous):

l is 45

OpenStudy (dumbcow):

no (1/2)*6*L = 45 solve for L by multiplying 6 and 1/2, then dividing on both sides

OpenStudy (anonymous):

i got 15

OpenStudy (dumbcow):

correct, good job

OpenStudy (anonymous):

The volume of a cylinder is 628 cm3. Find the radius of the base if the cylinder has a height of 8 cm. Round your answer to the nearest 0.1 cm.

OpenStudy (dumbcow):

ok for this one can you tell me the volume formula for cylinder?

OpenStudy (anonymous):

basexheight

OpenStudy (dumbcow):

good base is a circle, whats the area of a circle?

OpenStudy (anonymous):

pir^2

OpenStudy (dumbcow):

good V = (pir^2)*h pi*r^2*8 = 628

OpenStudy (dumbcow):

multiply 8*pi divide on both sides

OpenStudy (anonymous):

78.5

OpenStudy (dumbcow):

dont forget about pi (use3.14) r^2 = 78.5/pi divide 78.5/3.14 then take the square root to get r

OpenStudy (anonymous):

4.99

OpenStudy (dumbcow):

correct

OpenStudy (dumbcow):

actually with rounding to nearest .1 it would just be 5

OpenStudy (anonymous):

Sylvia has just discovered that the valve on her cement truck failed during the night and that all the contents ran out to form a giant cone of hardened cement. To make an insurance claim, she needs to figure out how much cement is in the cone. The circumference of its base is 44 feet, and it is 5 feet high. Calculate the volume to the nearest cubic foot.

OpenStudy (dumbcow):

to answer this we need to find radius what is circumference of circle

OpenStudy (anonymous):

so that mean 22

OpenStudy (dumbcow):

good but we still need to divide by pi (3.14)

OpenStudy (dumbcow):

2pi*r =44 pi*r=22 r=22/pi

OpenStudy (anonymous):

7.00

OpenStudy (dumbcow):

good next we use volume of cone formula

OpenStudy (anonymous):

so that i use 1/3bh

OpenStudy (anonymous):

really all you need to do to solve these problems is just write out the sets of equations that apply to the problem for example the circumference equation and the volume of a cone. then, you plug in what you know such as the measurement of circumference to solve for the radius then plug that into the volume of a cone formula.

OpenStudy (dumbcow):

yes b = area of circle w/ radius of 7

OpenStudy (anonymous):

use what you do know to find what you don't

OpenStudy (anonymous):

49pi

OpenStudy (dumbcow):

good V = 1/3 * 49pi*5 multiply it out

OpenStudy (anonymous):

with pi?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

256.56

OpenStudy (dumbcow):

good now round to nearest foot

OpenStudy (anonymous):

257

OpenStudy (anonymous):

don't forget your units

OpenStudy (dumbcow):

yep ok i gotta go keep doing these steps, find correct formula plug in numbers you you solve for ones you don't good luck

OpenStudy (dumbcow):

thanks kristo, forgot about that

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

A sealed rectangular container 6 cm by 12 cm by 15 cm is sitting on its smallest face. It is filled with water up to 5 cm from the top. How many centimeters from the bottom will the water level reach if the container is placed on its largest face?

OpenStudy (anonymous):

alright first you need to find the volume of the water on its smallest face

OpenStudy (anonymous):

1080

OpenStudy (anonymous):

one sec i will be right back...

OpenStudy (anonymous):

ok i'm back

OpenStudy (anonymous):

so what might be a good idea is to draw a picture first which face would be the smallest?

OpenStudy (anonymous):

i drew what next

OpenStudy (anonymous):

okay so what you have is a rectangular prism with a base of 6x12 (the smallest face) making the height without water 15

OpenStudy (anonymous):

so the height of the water is five cm from the top so you subtract five from fifteen to get the height of the water

OpenStudy (anonymous):

that make 72

OpenStudy (anonymous):

720 6x12x10 remember to include WxLxH

OpenStudy (anonymous):

then

OpenStudy (anonymous):

now we have the volume what we don't know is what the height of the water is when the container is set on its largest face

OpenStudy (anonymous):

so i would draw a picture of the new rectangle first what is the W and the L of the base?

OpenStudy (anonymous):

hint: largest face

OpenStudy (anonymous):

12 and 10

OpenStudy (anonymous):

12 and 15

OpenStudy (anonymous):

180

OpenStudy (anonymous):

we only subtracted five to find the height of the water when the container was on the small face, since we have the W and L now we solve for the height using the volume formula V=WxLxH

OpenStudy (anonymous):

use the volume formula we found before 720

OpenStudy (anonymous):

720=15x12xH

OpenStudy (anonymous):

i got 4 for height

OpenStudy (anonymous):

great

OpenStudy (anonymous):

now subtract that from six since it wants the height from the top of the prism

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you really understand how we got that?

OpenStudy (anonymous):

well some of the time

OpenStudy (anonymous):

try to image a cereal box, you have the box standing up. you know how high the cereal is from the top of the box you use that to find the volume of the cereal. then you tip the box on its larger face then you want to know what the height of the cereal is from the new top edge of the box

OpenStudy (anonymous):

does that help to picture the problem a little better?

OpenStudy (anonymous):

yes so BxH

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

or length times width times height = volume

OpenStudy (anonymous):

but you got the concept

OpenStudy (anonymous):

i have to sleep now but good luck

OpenStudy (anonymous):

we know that 4 is the height so we need to find the base